OpenStudy (anonymous):

Solve for y. a=x+y+z/3 y = a - 3x - z y = a - x - 3z y = 3a - x - z

Parth (parthkohli):

\(\Large \color{purple}{3a = x + y + z }\) \(\Large \color{purple}{ y = 3a - x - z }\) (By transposing)

OpenStudy (diyadiya):

\[a =x+y+ \frac{z}{3}\] Multiply each term by 3 Tell me what do you get?

Parth (parthkohli):

@Diyadiya It's the whole term divided by 3 :)